OpenStudy (anonymous):

What is the area of a circle with a diameter of twenty-eight meters? Use twenty-two sevenths for pi. six hundred sixteen square meters six hundred thirty-eight square meters six hundred sixty square meters six hundred eighty-two square meters @dg2

OpenStudy (anonymous):

@hello_j @xanny6

OpenStudy (anonymous):

area is pi r^2, so cut the diameter in half

OpenStudy (anonymous):

r=28/2=14

OpenStudy (anonymous):

put in area =pi*r^2 =3.14*14*14

OpenStudy (anonymous):

so area=615.44

OpenStudy (anonymous):

I got 615.44

OpenStudy (anonymous):

if use pi=22/7 then =616